OpenStudy (anonymous):

The area of a football field is x2 + 3x – 10. Find the length and width of the field in terms of x.

OpenStudy (anonymous):

area= LxW=(x+5)(x-2) so let's say L =(x_5) and W=(x-2)

OpenStudy (anonymous):

I mean (x+5)